OpenStudy (anonymous):

Find c if a = 16 and b = 30.

OpenStudy (anonymous):

a.c = 17 b.c = 1,156 c.c = 34 d.c = 68

OpenStudy (anonymous):

is there an equation for c, because otherwise this doesn't make sense.

OpenStudy (anonymous):

@tomo All im given is this: Say that a and b are the lengths of the legs of a right triangle and c is the length of the hypotenuse. Find c if a = 16 and b = 30.

OpenStudy (anonymous):

I mean its Pythagorean Theorem I just don't know how to set it up

OpenStudy (anonymous):

ok, since it is a right triangle you can use Pythagorean theorem. given a right triangle, where a and b are the two side lengths and c is the length of the hypotenuse, the following equation is true. a^2+b^2 = c^2 c = sqrt(a^2+b^2) solve for c.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw